OpenStudy (anonymous):

Part A: The sun produces 3.9 ⋅ 10^33 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 2.45 ⋅ 10^5 seconds?

OpenStudy (paxpolaris):



OpenStudy (anonymous):

can u explain?

OpenStudy (paxpolaris):

you need to multiply the number of "ergs per second" with the number of "seconds" to get ergs

OpenStudy (anonymous):

then i add the exponents or subtract or dont touch it?

OpenStudy (paxpolaris):

\[\Large 3.9\cdot 10^{33} \times 2.45 \cdot 10^5=9.555\cdot10^{33+5}\]

OpenStudy (anonymous):

thank u